Walt made an extra $9000.00 lst year from his part time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770.00 in interest. How much was invested at 8%?
Answer by checkley77(12569)

.09x+.08(9,000-x)=770
.o9x+720-.08x=770
.01x=770-720
.01x=50
x=50/.01
x=5,000 invested @ 9%
9,000-5,000=4,000 invested @ 8%
proof:
.09*5000+.08*4000=770
450+320=770
770=770